# A Human-Centered Approach to Identifying Promises, Risks, & Challenges of Text-to-Image Generative AI in Radiology

### v1

@@ Introduction
Traditionally, synthetic image generation has been used to address pivotal technical medical imaging tasks (Sorin et al. 2020), such as image reconstruction or correction (Levac et al. 2024). Now, with foundation T2I models becoming widely accessible, the allure of simulating rare conditions, creating patient vignettes, or augmenting training datasets is attracting significant interest (Thesen, Alilonu, and Stone 2024; Pérez-García et al. 2025; Waikel et al. 2023; Waisberg et al. 2023). However, much of this enthusiasm is driven by technological evaluations (Noel 2024; Bhardwaj et al. 2023; Lin et al. 2024; Ku et al. 2023) rather than human-centered ones, leading to a disconnect from key stakeholders’ real needs and challenges. As recently emphasized by human-centered AI researchers, evaluations need to extend beyond algorithmic metrics by understanding how the AI can address stakeholders’ real needs and what challenges arise by doing so (Sivaraman et al. 2025). Although recent research captures medical stakeholders’ perspectives on GenAI’s potential applications when used in practice (Yildirim et al. 2024), few works capture medical stakeholders’ perspectives on the technical challenges and domain-specific risks of generating synthetic medical imagery.
@@
Our formative discussions included six radiologists and one senior trainee, and they took place at the end of the quantitative evaluation of the model in the user study conducted in (Xu et al. 2024). These discussions focused on capturing stakeholders’ ideas on how the model might be applied in practice (RQ1). This informed the focus of tasks we gave participants during the second phase (model exploration), broadened our network for recruiting additional participants, and informed the selection of participant demographics. For the second phase, we recruited two medical students, three radiology trainees, and four radiologists (including two from the formative discussions) to prompt and evaluate the model outputs in a semi-structured interview format. Participants interacted with the model through an open-source medical imaging interface (created by Ziegler et al. (2020)) that we integrated the text-to-CT Scan model into for easy prompting and output review. By interacting with the model, participants imagined scenarios of how it can address real needs and challenges faced throughout radiology workflows (RQ1). The exploratory prompting activity also shed light on the technical challenges and potential domain-specific risks that researchers should consider when developing medical T2I models (RQ2).
Our work makes three contributions to the AIES and broader Responsible AI (RAI) community. First, we present a text-to-CT Scan GenAI plugin for a popular open-source medical imaging viewer (created by Ziegler et al. (2020)), which researchers can extend to explore additional approaches for human-centered evaluations of the ethical and safety challenges of domain-specific T2I models. Second, we are the first work to leverage a human-centered approach to explore a medical T2I model with medical stakeholders. As a result of this, our paper expands upon Yildirim et al. (2024)’s human-centered GenAI work for applications of multimodal AI in radiology by mapping out applications of domain-specific T2I across medical education, radiology training, and practice. Third, we extend existing RAI work by presenting technical challenges and domain-specific risks that emerged from participants interacting with the model, extending del Cerro and Esteve (2023)’s position on the implications of using medical T2I GenAI. These challenges and risks range from topics such as confirmation bias, misrepresentations of image findings, and output image resolution. We discuss the implications of developing medical T2I and suggest future research directions to consider exploring.
@@
Several studies have recently been evaluating foundation models’ capabilities to generate medical images (Bhardwaj et al. 2023; Moin et al. 2024; Temsah et al. 2024)Noel (2024) and Kumar et al. (2024) similarly probe different text-to-image GenAI models to evaluate their capability in generating different medical images. One study ran an educational intervention study with students to quantitatively evaluate the impact of generative AI images in education about pediatric genetics (Waikel et al. 2023)Thesen, Alilonu, and Stone (2024) recruited a diverse group of stakeholders to evaluate generated images for education across several scales, including the images’ potential to be used in medical presentations. The above-mentioned studies primarily focus on anatomical realism or the presence of diseases prompted by the research team instead of assessing whether T2I capabilities meaningfully address stakeholders’ needs and challenges. Our work addresses this gap by having stakeholders evaluate the promises, risks, and challenges of T2I capabilities in the context of medical education, training, and practice.
@@
Numerous previous works from medical literature have explored foundation models’ ability to generate images sufficient to use for medical education (Waikel et al. 2023; Kumar et al. 2024; Koljonen 2023; Zhu et al. 2024; Noel 2024). For example,  Zhu et al. (2024) probe the potential of OpenAI’s DALL-E (Ramesh et al. 2021) to generate ECGs, including ECGS that are abnormal for teaching purposes. Although they do not present a study on probing or interviewing stakeholders, the authors use their professional experiences to provide insight into DALL-E’s potential as a teaching tool. Currie, Hawk, and Rohren (2024) raises numerous concerns about using T2I models in nuclear medicine for education, citing the need to identify policies and regulations that address misuse and abuse, misrepresentations and hallucinations, and patient privacy concerns.
In Clinical Practice. Clinical decision-making informed by medical imaging is a nuanced process that relies on expertise, peer collaboration, iterative review, and the use of reference resources to augment diagnostic accuracy and patient care. Despite advancements in AI, existing tools and interfaces tend to overlook these needs and focus on higher-order problems such as pathology detection (Sloan et al. 2024) or automated report writing (Yousefirizi et al. 2022). However, recent work has argued for the need to explore AI’s true value to clinicians throughout their clinical workflows (Sivaraman et al. 2025). A recent study by Waisberg et al. (2023) look at text-to-image GenAI’s potential as a collaborative tool for clinicians in helping them understand neuro-visual diseases’ impact on patients’ ability to see by generating an image based on a description of the patient’s symptoms. However, no formal quantitative or human-centered evaluations were conducted. del Cerro and Esteve (2023) suggests how T2I can be used by physicians to simulate treatment effects to navigate treatment plans or by radiologists to enhance noisy images. The authors simultaneously cite general concerns and potential implications of misuse and data bias.
Although previous works have identified risks with T2I models in medical education, training, and practice, they are based on domain-agnostic T2I models, such as DALL-E or MidJourney. Furthermore, they primarily consist of showing already generated images to participants instead of allowing key stakeholders to explore and prompt the model themselves. Our work builds on the existing literature by having medical students, radiology residents, and radiologists interact with a domain-specific T2I to capture richer applications and challenges of medical T2I.
@@
Figure 2: Open-sourced medical imaging interface (Ziegler et al. 2020) with our GenAI plugin powered by MedSyn (Xu et al. 2024). (2) Patient Case. The first column shows a real patient’s case consistent with severe right pleural effusion (sampled from CT-RATE dataset (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024). (3) CT Scan Viewing Window. The second and third columns show a CT scan comparison viewing window, where a human’s CT scan (from (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024)) is on the left, and a generated CT scan from MedSyn (Xu et al. 2024) is on the right. (1) Prompt UI. The fourth column is our text-to-CT Scan GenAI plugin, allowing users to generate a CT scan.
@@
We interviewed six radiologists and one senior radiology trainee (F1-F7) for an average of $6.86.8$ minutes after they quantitatively evaluated outputs from the MedSyn model [^1] Additional details and results in supplemental materials.. Ultimately, the formative discussions provided us with preliminary insights into radiologists’ challenges and their perceptions of text-to-CT Scan GenAI. Participants imagined how T2I could augment education, training, and practice, but struggled to detail specific scenarios since many had never used a T2I model before. We also uncovered radiologists’ challenges with current online resources, such as visualizing the exact pathologies they search for, the importance of visual comparative analyses, and their lack of exposure to T2I GenAI. These challenges inspire our selection of a user interface and design choices made with our GenAI plugin.
@@
Our formative discussions with radiologists revealed to us that they frequently search for image findings and pathologies on domain-specific search engines. However, they may not always find exactly what they are searching for. Thus, we developed a GenAI plugin for an open-source medical image viewer (Ziegler et al. 2020),  [^2] Publicly accessible demo is available here: https://genai-radiology.web.app/login. Please note that the model server to generate images is only turned on for interviews. powered by a text-to-CT Scan GenAI model, MedSyn (Xu et al. 2024). The MedSyn model enables users to generate CT scans based on exact and specific pathologies, extending capabilities beyond current search tools. As depicted in Figure 1, MedSyn takes a user’s text prompt and generates a 3D CT Scan consisting of the axial, sagittal, and coronal views of the lungs. Generating the complete CT scan takes between three and four minutes when using an A100 GPU. Our plugin manifests in the interface (shown in Figure 2) as a text box where users can (1a) name their prompt, (1b) write their prompt, and (1c) generate a CT scan by calling the model API through the click of a button. Users can view a list of previous CT scans that they generated (1d) and have the opportunity to provide feedback on the generation (1e).
@@
When using this tool, a radiologist can view a patient’s case history (2a), noting their findings (2b) and impressions (2c). This functionality is incorporated to align with how radiologists currently read CT scans in clinical practice. Viewing the patient’s CT in the main viewer (3a) gives them access to numerous functionalities (3b) that they are used to having, such as adjusting the contrast to preset window levels (e.g., soft tissue, bone, and lung) [^3] Presets were determined by an author, who is a radiologist., adding annotations, comparing images side-by-side, and using a localizer tool. Many of these functionalities expand upon what the existing radiology search engines offer. For example, they allow radiologists to view and compare generated CT scans with real patient data as well as manipulate and annotate the images.
@@
We conducted 45-minute semi-structured interviews with relevant medical stakeholders to elicit their perspectives on the potential applications and implications of text-to-CT Scan GenAI. These interviews were designed to be exploratory in nature; the goal was not to formally audit or analyze prompting behaviors. The interview sessions consist of four parts, inspired by Cai et al. (2019)’s approach with a clinical decision-making design probe, which included a pre-probe, probe, and post-probe phase. We describe each part below.
@@ Introduction.
Participants were asked about their experiences with text-to-image GenAI. For participants who have never used text-to-image GenAI before, we showed them a short conversation made in ChatGPT where we already requested the generation of a medical image from DALL-E [^4] Provided in the supplemental materials.. All participants were asked before interacting with the model if they could think of any use cases for text-to-image GenAI being valuable to them throughout their education, training, or practice to prime them for the rest of the interview session. Following this, we showed all participants an animated illustration of how the text-to-CT Scan GenAI model works, and participants were acclimated to the interface through a guided walkthrough video before interacting with the interface themselves.
@@
During the probe phase, participants generated two to four CT scans, depending on session duration and the depth of discussions. They were guided to use the model to generate CT scans based on different goals, such as validating their impression of a real patient’s CT scan loaded in the viewer from CT-RATE (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024), learning about a pathology through differential diagnosis, and generating something of their choice. Throughout this process, participants provided feedback about the prompts they chose, the quality of the generation, and insights on the model’s potential applications in radiology education, training, and practice.
@@
Participants were recruited through snowball sampling methods from our clinical collaborators’ networks. For each session, we provided the participants with an overview of the study’s purpose and methods and obtained their consent for participation. All participants consented to audio and screen recording for data analysis and were provided a form of monetary compensation for their time. We interviewed nine stakeholders (R1–R9) across three different institutions: two medical students, three radiology trainees, and four attending radiologists (two from the formative study and two new ones); experience distribution is provided in Table 1). Sessions were held on Zoom or Teams, based on preference, and lasted for an average of $46.3446.34$ minutes.
@@
We used an affinity diagramming approach to analyze the qualitative data. First, the transcripts from the interview sessions were verified for accuracy, and corrections were made by listening to the recordings as necessary. Ideas from the transcripts were pulled out as quotes, and affinity mapping was conducted to group similar ideas across interviews based on common themes. Multiple authors sequentially reviewed the groupings individually, and all the quotes were grouped into primary themes that aligned with our research questions. Sub-themes were collectively generated and agreed upon after the quotes’ groups were finalized.
@@
We first report on the identified promises of T2I in medical education, training, and practice (RQ1). We then follow up on these promises with the risks and challenges of T2I that stakeholders identified (RQ2). Table 2 and Table 3 show high-level representations of the findings for both research questions. A complete list of the prompts created by each participant, along with their comments about the prompt and the generated output, is provided in the supplemental materials.
@@
An application context that was brought up frequently in the formative discussions, as well as our semi-structured interviews, was the potential to use this technology in medical education and training. R7 presents a compelling case for using text-to-CT Scan GenAI in medical education because, “…it gives you the opportunity to identify and recognize something that is potentially super dangerous, that you then don’t have to guess about when you see it in real life”. We see the essence of this quote throughout our conversations with students and trainees. We have grouped participants’ insights on specific use cases into three themes regarding the promises of T2I GenAI in medical education and training: (1) supplementing traditional lectures and practice materials, (2) expediting learning through variations, differentials, and rare cases, and (3) empowering trainees during on-call or emergency situations.
@@
Despite numerous resources out there, “…there isn’t a great place to be like, Hey med students, let’s go here and learn chest CT today” (R8). R8 goes on further describing how they have never done anything in their classes where they get to interact with and compare the differences between abnormal and normal anatomies on CT scans. This could be due to how medical school lectures are designed, showing static images on lecture slides. R3 corroborates this limitation in existing lecture style by describing how they get content for their lecture slides by searching online or through their archives: “I have this list of typical cases for all of these different lesions…choosing representative images from them…or showing a student the entirety of the scan”. However, R3 went on to envision a scenario where it could generate specific examples to use on their slides: “I’m thinking about pre-populating…I’m gonna produce a lecture on this list of obscure topics…And I need specific examples of A, B, C, … I feed those into your generator, and…I’ll come back in an hour….and I’m ready to start producing my lecture.” R1 provides another perspective, acknowledging how convenient text-to-CT Scan GenAI could be when “…teaching the residents or medical students, and we just want to show them something, we can just type it, and it will pop up”.
@@
As medical students transition into their training program, it becomes increasingly important to get as much exposure as possible to a wide variety of cases. As R1 says, “the more you see, the better radiologist you will become”. However, despite relying on existing resources populated by the community and their training to see as many diverse images as possible, they are limited by what others can share based on patient privacy regulations or what patient cases come through at their hospital. Radiology trainee, R2, highlighted how this GenAI could accelerate their learning by allowing them to study variations of diseases, “this would expedite that if I could just be like, ‘okay, I need to know the difference between this, this, and this’…I can just generate 5 different variances of those diseases on this tool with the different manifestations of those variants.
@@
Beyond variations and differentials, R8 imagined how this GenAI could be applied in day-to-day training scenarios to prepare for region-specific diseases. For example, they described a scenario where generated CT scans could be “…included in my docket of cases I have to look at for the day and one or two of them might be training, region-specific training”. They similarly mentioned the tool’s value in training medical professionals in low-resource environments, such as being a “…benefit to even Africa…And being able to say, ‘Okay, here’s what Cardiomegaly looks like with someone who might have tuberculosis”’.
@@
R2 expands this idea by imagining how they could use the tool to generate baseline CT scans to make sure they do not miss anything while on-call. R7 imagines how generating images for severe conditions ahead of shifts in the ER could accelerate trainees’ decision-making, empowering them to handle emergency situations appropriately and promptly: “…if you can generate your own images of what different aortic dissections look like, then, when you have a patient that comes into the ER with ripping chest pain, and you take them to CT, and you see, like, ‘Oh, sh*t! This is really bad, I need to call someone for help now’, as opposed to ‘Okay, I haven’t seen this before, but this sounds like what I think it could be like. Let me wait and see if I can get CT to read this, and then I have to call my attending’, and that all adds time in an emergency situation…time is life”. R7 then prompted the model to generate an aortic dissection.
@@
The comparative functionality in the interface allowed R6 to discover that their impression of the original patient case presented to them was not completely correct, when they compared the image findings in the generated image to the provided real CT: “So, I said, smaller effusion, and sometimes people say mild effusion, some say severe effusion. But so this generated AI gives me a very small effusion. But now I’m thinking this might be a mild effusion instead of a small effusion, which can be very helpful actually…Something like when we use those subjective words and if we have something where somethings objective, could help us a lot actually”.
Together, these scenarios underscore the potential of text-to-CT Scan GenAI to enhance the accuracy, efficiency, and precision of image interpretation in clinical practice, empowering clinicians with both routine and challenging cases. However, it also raises implications for accountability, confirmation bias, and unintentional deskilling, which we will discuss in Section Risks & Challenges with Medical T2I GenAI.
@@
An interesting use case outside of direct diagnosis or education that was mentioned was in planning and communicating with other stakeholders. For example, R2 described how generating CT scans could be “super helpful in surgical planning…you can show the patient like, ‘hey, this is going to be what you look like post-surgery.” In similar vain, R8 describes how they could imagine a scenario showing patients how their condition is looking.
As previously mentioned, radiologists are not the only ones who will be interacting with medical imaging throughout the clinical workflow. This could introduce multiple impressions that may disagree about a diagnosis. As such, R5 brings up how “These generated images can act as additional evidence when I am communicating my diagnosis to other doctors”. R7 motivates R5’s point, noting how some surgeons they have shadowed will “…trust their own reading of the image over whatever the official [radiologist’s] read is”. This is an interesting dynamic in the clinical workflow where generated CT scans could aid in these collaborative workflows, helping radiologists provide validations of their reports by generating CTs that show differentials for other physicians to read.

### v2

@@ Introduction
Traditionally, synthetic image generation has been used to address pivotal technical medical imaging tasks (Sorin et al. 2020), such as image reconstruction or correction (Levac et al. 2024). Now, with foundation T2I models becoming widely accessible, the allure of simulating rare conditions, creating patient vignettes, or augmenting training datasets is attracting significant interest (Thesen, Alilonu, and Stone 2024; Pérez-Garc´ıa et al. 2025; Waikel et al. 2023; Waisberg et al. 2023). However, much of this enthusiasm is driven by technological evaluations (Noel 2024; Bhardwaj et al. 2023; Lin et al. 2024; Ku et al. 2023) rather than human-centered ones, leading to a disconnect from key stakeholders’ real needs and challenges. As recently emphasized by human-centered AI researchers, evaluations need to extend beyond algorithmic metrics by understanding how the AI can address stakeholders’ real needs and what challenges arise by doing so (Sivaraman et al. 2025). Although recent research captures medical stakeholders’ perspectives on GenAI’s potential applications when used in practice (Yildirim et al. 2024), few works capture medical stakeholders’ perspectives on the technical challenges and domain-specific risks of generating synthetic medical imagery.
@@
Our formative discussions included six radiologists and one senior trainee, and they took place at the end of the quantitative evaluation of the model in the study conducted by (Xu et al. 2024). These discussions focused on capturing stakeholders’ ideas on how the model might be applied in practice (RQ1). This informed the type of tasks we gave participants during the second phase (model exploration), broadened our network for recruiting additional participants, and informed the selection of participant demographics. For the second phase, we recruited two medical students, three radiology trainees, and four radiologists (including two from the formative discussions) to prompt and evaluate the model in a semi-structured interview format. To facilitate easy prompting and output review, participants interacted with the model through a plugin that we developed for an existing open-source medical imaging interface (created by Ziegler et al. (2020)). By interacting with the model, participants imagined scenarios of how it can address real needs and challenges faced throughout radiology workflows (RQ1). The exploratory prompting activity also shed light on the technical challenges and potential domain-specific risks that researchers should consider when developing medical T2I models (RQ2).
Our work makes three contributions to the AIES and broader Responsible AI (RAI) community. First, we present a text-to-CT Scan GenAI plugin for a popular open-source medical imaging viewer (created by Ziegler et al. (2020)), which researchers can extend to explore additional approaches for human-centered evaluations of the ethical and safety challenges of domain-specific T2I models. Second, we are the first work to leverage a human-centered approach to explore a medical T2I model with medical stakeholders. As a result of this, our paper expands upon Yildirim et al. (2024)’s human-centered GenAI work for applications of multimodal AI in radiology by mapping out applications of domain-specific T2I across medical education, radiology training, and practice. Third, we extend existing RAI work and build on del Cerro and Esteve (2023)’s position on the implications of using medical T2I GenAI by presenting technical challenges and domain-specific risks that emerged from participants interacting with the model. These challenges and risks range from topics such as confirmation bias, misrepresentation of image findings, and output image resolution. We discuss the implications of developing medical T2I and suggest future research directions to consider exploring.
@@
Several studies have recently evaluated the capability of domain-agnostic foundation models to generate medical images (Bhardwaj et al. 2023; Moin et al. 2024; Temsah et al. 2024)Noel (2024) and Kumar et al. (2024) similarly probe different text-to-image GenAI models to evaluate their capability in generating different medical images. One study ran an educational intervention study with students to quantitatively evaluate the impact of generative AI images in education about pediatric genetics (Waikel et al. 2023)Thesen, Alilonu, and Stone (2024) recruited a diverse group of stakeholders to evaluate generated images for education across several scales, including the images’ potential to be used in medical presentations. The above-mentioned studies primarily focus on quantitative evaluations of domain-agnostic T2I models with proxy metrics, such as anatomical realism or the presence of diseases, instead of assessing whether T2I capabilities meaningfully address stakeholders’ needs and challenges. Our work builds on existing studies by uncovering the promises, risks, and challenges of T2I capabilities in the context of medical education, training, and practice through prompting activities with relevant stakeholders.
@@
Numerous previous works from medical literature have explored foundation models’ ability to generate images sufficient to use for medical education (Waikel et al. 2023; Kumar et al. 2024; Koljonen 2023; Zhu et al. 2024; Noel 2024). For example, Zhu et al. (2024) probe the potential of OpenAI’s DALL-E (Ramesh et al. 2021) to generate ECGs, including ECGS that are abnormal for teaching purposes. Although they do not present a study on probing or interviewing stakeholders, the authors use their professional experiences to provide insight into DALL-E’s potential as a teaching tool. Currie, Hawk, and Rohren (2024) raise numerous concerns about using T2I models in nuclear medicine for education, citing the need to identify policies and regulations that address misuse and abuse, misrepresentations and hallucinations, and patient privacy concerns. Our paper adds to these findings by capturing perspectives on T2I in medical education from a diverse group of medical stakeholders.
In Clinical Practice. Clinical decision-making informed by medical imaging is a nuanced process that relies on expertise, peer collaboration, iterative review, and the use of reference resources to augment diagnostic accuracy and patient care. Despite advancements in AI, existing tools and interfaces tend to overlook these needs and focus on higher-order problems such as pathology detection (Sloan et al. 2024) or automated report writing (Yousefirizi et al. 2022). However, recent work has argued for the need to explore AI’s true value to clinicians throughout their clinical workflows (Sivaraman et al. 2025). A recent study by Waisberg et al. (2023) looks at text-to-image GenAI’s potential as a collaborative tool for clinicians in helping them understand neuro-visual diseases’ impact on patients’ ability to see by generating an image based on a description of the patient’s symptoms. However, no formal quantitative or human-centered evaluations were conducted. del Cerro and Esteve (2023) suggests how T2I can be used by physicians to simulate treatment effects to navigate treatment plans or by radiologists to enhance noisy images. The authors simultaneously cite general concerns and potential implications of misuse and data bias.
Although previous works have identified risks with T2I models in medical education, training, and practice, they are primarily based on domain-agnostic T2I models, such as DALL-E or MidJourney. Furthermore, they primarily consist of showing already generated images to participants instead of allowing key stakeholders to explore and prompt the model themselves. Our work builds on the existing literature by having medical students, radiology residents, and radiologists interact with a domain-specific T2I to capture richer applications and challenges of medical T2I.
@@
Figure 2: Open-sourced medical imaging interface by Ziegler et al. (2020) with our GenAI plugin powered by MedSyn (Xu et al. 2024). (2) Patient Case. The first column shows a real case consistent with severe right pleural effusion (sampled from CT-RATE dataset (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024). (3) CT Scan Viewing Window. The second and third columns show a CT scan comparison viewing window, where a human’s CT scan (from (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024)) is on the left, and a generated CT scan from MedSyn (Xu et al. 2024) is on the right. (1) Prompt UI. The fourth column is our text-to-CT Scan GenAI plugin, allowing users to generate a CT scan.
@@
We interviewed six radiologists and one senior radiology trainee (F1-F7) for an average of $6.8$ minutes after they quantitatively evaluated outputs from the MedSyn model in Xu et al. (2024)’s study [^1] Provided in Appendix Section Formative Study. Ultimately, the formative discussions provided us with preliminary insights into radiologists’ challenges and their perceptions of the value of text-to-CT Scan GenAI in education, training, and practice. However, participants struggled to detail specific scenarios since many had never used T2I GenAI before. We also uncovered participants’ challenges with current online resources, such as visualizing the exact pathologies they search for, the importance of visual comparative analyses, and lack of exposure to T2I GenAI. These challenges inspire us to integrate MedSyn into the medical imaging interface by Ziegler et al. (2020).
@@
Our formative discussions with radiologists revealed to us that they frequently search for image findings and pathologies on domain-specific search engines. However, they may not always find exactly what they are searching for. Therefore, we developed a GenAI plugin powered by the text-to-CT Scan GenAI model, MedSyn (Xu et al. 2024), and integrated it into a popular open-source medical image viewer (Ziegler et al. 2020) [^2] Plugin Demo available at https://genai-radiology.web.app/.. The MedSyn model enables users to generate CT scans based on exact and specific pathologies, extending capabilities beyond current search tools. As depicted in Figure 1 below, MedSyn takes a user’s text prompt and generates a 3D CT Scan consisting of the axial, sagittal, and coronal views of the lungs. Generating the complete CT scan takes between three and four minutes when using an A100 GPU. Our plugin manifests in the interface (shown in Figure 2 below) as a text box where users can (1a) name their prompt, (1b) write their prompt, and (1c) generate a CT scan by calling the model API through the click of a button. Users can view a list of previous CT scans that they generated (1d) and have the opportunity to provide feedback on the generation (1e).
@@
When using this tool, a radiologist can view a patient’s case history (2a), noting their findings (2b) and impressions (2c). This functionality is incorporated to align with how radiologists currently read CT scans in clinical practice. Viewing the patient’s CT in the main viewer (3a) gives them access to numerous functionalities (3b) that they are used to having, such as adjusting the contrast to preset window levels (e.g., soft tissue, bone, and lung) [^3] Presets were determined by an author, who is a radiologist., adding annotations, comparing images side-by-side, and using a localizer tool. Many of these functionalities either expand upon what the existing radiology search engines offer or leverage findings from our formative interviews. For example, the tool allows users to view and compare a generated CT scan with a real CT scan, as well as annotate them.
@@
We conducted 45-minute semi-structured interviews with relevant medical stakeholders to elicit their perspectives on the potential applications and implications of text-to-CT Scan GenAI. These interviews were designed to be exploratory in nature; the goal was not to formally audit or analyze prompting behaviors. The interview sessions consist of four parts, inspired by Cai et al. (2019)’s approach with a clinical decision-making design probe, which included a pre-probe, probe, and post-probe phase. We describe each part in the following subsections.
@@ Introduction.
Participants were asked about their experiences with text-to-image GenAI. For participants who have never used text-to-image GenAI before, we showed them a short conversation made in ChatGPT where we already requested the generation of a medical image from DALL-E [^4] Provided in Appendix Section Pre-Probe Materials.. All participants were asked before interacting with the model if they could think of any use cases for text-to-image GenAI being valuable to them throughout their education, training, or practice to prime them for the rest of the interview session. Following this, we showed all participants an animated illustration of how the text-to-CT Scan GenAI model works, and participants were acclimated to the interface through a guided walkthrough video.
@@
During the probe phase, participants generated two to four CT scans, depending on session duration and the depth of discussions. They were guided to use the model to generate CT scans based on different goals informed by the formative interviews, such as validating their impression of a real patient’s CT scan loaded in the viewer from CT-RATE (Hamamci et al. 2023, 2024; Hamamci, Er, and Menze 2024), learning about a pathology through differential diagnosis, and generating something of their choice. Throughout this process, participants provided feedback about the prompts they chose, the quality of the generation, and insights on the model’s potential applications in radiology education, training, and practice.
@@
Participants were recruited through snowball sampling methods from our clinical collaborators’ networks. For each session, we provided the participants with an overview of the study’s purpose and methods and obtained their consent for participation. All participants consented to audio and screen recording for data analysis and were provided a form of monetary compensation for their time. We interviewed nine stakeholders (R1–R9) across three different institutions: two medical students, three radiology trainees, and four attending radiologists (two from the formative study and two new ones); experience distribution is provided in Table 1). Sessions were held on Zoom or Teams, based on preference, and lasted for an average of $46.34$ minutes.
@@
We used an affinity diagramming approach to analyze the qualitative data. First, the transcripts from the interview sessions were verified for accuracy, and corrections were made by listening to the recordings as necessary. Quotes were pulled out from the transcripts, and affinity mapping was conducted by multiple authors to group similar quotes across interviews based on common themes. All quotes were grouped into primary themes and sub-themes that aligned with our research questions. Multiple authors grouped the quotes individually, and then the final groups were determined through discussion.
@@
We first report on the identified promises of T2I in medical education, training, and practice (RQ1). We then follow up on these promises with the risks and challenges of T2I GenAI that stakeholders identified (RQ2). Table 2 and Table 3 show high-level representations of the findings for both research questions. A complete list of the prompts created by each participant, along with their comments about the prompt and the generated output, is provided in the supplemental materials.
@@
An application context that was brought up frequently in the formative interviews, as well as our semi-structured interviews, was the potential to use this technology in medical education and training. R7 presents a compelling case for using text-to-CT Scan GenAI in medical education because, “…it gives you the opportunity to identify and recognize something that is potentially super dangerous, that you then don’t have to guess about when you see it in real life”. We see the essence of this quote throughout our conversations with students and trainees. We have grouped participants’ insights on specific use cases into three themes regarding the promises of T2I GenAI in medical education and training: (1) supplementing traditional lectures and practice materials, (2) expediting learning through variations, differentials, and rare cases, and (3) empowering trainees during on-call or emergency situations.
@@
Despite numerous resources out there, “…there isn’t a great place to be like, Hey med students, let’s go here and learn chest CT today” (R8). R8 goes on further describing how they have never done anything in their classes where they get to interact with and compare the differences between abnormal and normal anatomies on CT scans. This could be due to how medical school lectures are designed, showing static images on lecture slides. R3 corroborates this limitation in existing lecture style by describing how they get content for their lecture slides by searching online or through their archives: “I have this list of typical cases for all of these different lesions…choosing representative images from them…or showing a student the entirety of the scan”. However, R3 went on to envision a scenario where they could use this T2I model to generate specific examples to use on their slides: “I’m thinking about pre-populating…I’m gonna produce a lecture on this list of obscure topics…And I need specific examples of A, B, C, … I feed those into your generator, and…I’ll come back in an hour….and I’m ready to start producing my lecture.” R1 provides another perspective, acknowledging how convenient text-to-CT Scan GenAI could be when “…teaching the residents or medical students,…we just want to show them something, we can just type it, and it will pop up.”
@@
As medical students transition into their training program, it becomes increasingly important to get as much exposure as possible to a wide variety of cases. As R1 says, “the more you see, the better radiologist you will become”. However, despite relying on existing resources populated by the community and their training to see as many diverse images as possible, they are limited by what others can share based on patient privacy regulations or what patient cases come through at their hospital. Radiology trainee, R2, highlighted how this GenAI could accelerate their learning by allowing them to study variations of diseases: “…this would expedite that if I could just be like, ‘okay, I need to know the difference between this, this, and this’…I can just generate 5 different variances of those diseases on this tool with the different manifestations of those variants.
@@
Beyond variations and differentials, R8 imagined how this GenAI could be applied in day-to-day training scenarios to prepare for region-specific diseases. For example, they described a scenario where generated CT scans could be “…included in my docket of cases I have to look at for the day and one or two of them might be training, region-specific training”. They similarly mentioned the tool’s value in training medical professionals in low-resource environments, such as being a “…benefit to even Africa…And being able to say, ‘Okay, here’s what Cardiomegaly looks like with someone who might have tuberculosis.”’
@@
R2 expands this idea by imagining how they could use the tool to generate baseline CT scans to make sure they do not miss anything while on-call. R7 imagines how generating images for severe conditions ahead of shifts in the ER could accelerate trainees’ decision-making, empowering them to handle emergency situations appropriately and promptly: “…if you can generate your own images of what different aortic dissections look like, then, when you have a patient that comes into the ER with ripping chest pain, and you take them to CT, and you see, like, ‘Oh, sh*t! This is really bad, I need to call someone for help now’, as opposed to ‘Okay, I haven’t seen this before, but this sounds like what I think it could be like. Let me wait and see if I can get CT to read this, and then I have to call my attending’, and that all adds time in an emergency situation…time is life.” R7 then prompted the model to generate an aortic dissection.
@@
The comparative functionality in the interface allowed R6 to discover that their impression of the original patient case presented to them was not completely correct when they compared the image findings in the generated image to the provided real CT: “So, I said, smaller effusion, and sometimes people say mild effusion, some say severe effusion. But so this generated AI gives me a very small effusion. But now I’m thinking this might be a mild effusion instead of a small effusion, which can be very helpful actually…Something like when we use those subjective words and if we have something where somethings objective, could help us a lot actually.
Together, these scenarios underscore the potential of text-to-CT Scan GenAI to enhance the accuracy, efficiency, and precision of image interpretation in clinical practice, empowering clinicians while reviewing both routine and challenging cases. However, it also raises implications for accountability, confirmation bias, and unintentional deskilling.
@@
An interesting use case outside of direct diagnosis or education that was mentioned was in planning and communicating with other stakeholders. For example, R2 described how generating CT scans could be “super helpful in surgical planning…you can show the patient like, ‘hey, this is going to be what you look like post-surgery.” In a similar vein, R8 describes how they could imagine a scenario where they show a patient how their condition looks compared to a normal condition.
Radiologists are not the only ones who will be interacting with medical imaging throughout the clinical workflow. This could introduce multiple impressions that may disagree about a diagnosis. As such, R5 brings up how “These generated images can act as additional evidence when I am communicating my diagnosis to other doctors.” R7 motivates R5’s point, noting how some surgeons they have shadowed will “…trust their own reading of the image over whatever the official [radiologist’s] read is.” This is an interesting dynamic in the clinical workflow where generated CT scans could aid in these collaborative workflows, helping radiologists provide validations of their reports by generating CTs that show differentials for other physicians to read.